Two small bodies of masses 2.00 kg and 4.00 kg are kept at rest at a separation of 2.0 m. Where should a particle of mass 0.10 kg be placed to experience no net gravitational force from these bodies? The particle is placed at this point. What is the gravitational potential energy of the system of three particles with usual reference level?

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. Let 0.1 kg man is x m from 2kg mass and (2 – x) m from 4 kg mass. ∴ 2 * 0.1/x^2 = - 4 * 0.1/ (2 - x)^2 Or 0.2/x^2 = - 0.4/(2 - x)^2 Or 1/x^2 = 2/(2 - x)^2 or (2 - x)^2 = 2 x^2 Or 2 – x = √2 x or x(r base 2 + 1)^2 = 2 Or x = 2/2.414 = 0.83 m from 2kg mass.